Billy's restaurant ordered 200 flowers for Mother's Day. They ordered carnations at $1.50 each, roses at $5.75 each, and daisies at $2.60 each. they ordered mostly carnations, and 20 fewer roses than daisies. the total order came to $589.50. How do i find out how many of each type of flower was ordered?
Answer by stanbon(75887)

Billy's restaurant ordered 200 flowers for Mother's Day. They ordered carnations at $1.50 each, roses at $5.75 each, and daisies at $2.60 each. they ordered mostly carnations, and 20 fewer roses than daisies. the total order came to $589.50. How do i find out how many of each type of flower was ordered?
------------------------
Quantity Eq: c + r + d = 200
Value Eq:::1.5c+5.75r+2.6d = 589.5 dollars
r = d-20
c = 200-d-(d-20) = 200-2d+20 = 220-2d
---
Three Equations:
1.5c+5.75r+2.6d = 589.5 dollars
c......0*r..+2d...= 220
c + r + d = 200
--------------------
I used a Matrix Method to get:
c = 80 (# of carnations-
r = 50 (# of roses)
d = 70 (# of daisies)
==========================
Cheers,
Stan H.
================